OpenStudy (anonymous):

Use the table below to find the slope . Months : 4 , 5 , 6 , 7 , 8 , 10 . Weight in pounds : 11, 16 , 21 , 26 , 31 , 41 . Explain your process for finding the slope of the table .

OpenStudy (anonymous):

make a graph use months as x values along bottom use weight as y values up left side

OpenStudy (anonymous):

But it not asking for that its asking to explain in words of how to find the slope . I already found the differences what do I do now

OpenStudy (anonymous):

the formula to determine the slope: y2 - y1 ------ = slope x2 - x1 Use the first and last ordered pairs (points on graph) (11,4) and (41,10) x1, y1 x2, y2

OpenStudy (anonymous):

(you needed to understand the graph in order to calculate slope)

OpenStudy (anonymous):

ok so i wrote down 10 - 4 _______ = 6/31 41-10 is this right so far ?

OpenStudy (anonymous):

yes, 6/31 is slope

OpenStudy (anonymous):

wait I put your values backwards

OpenStudy (anonymous):

should have been (4,11) and 10, 41)

OpenStudy (anonymous):

months need to be x values not y, my bad

OpenStudy (anonymous):

its ok (: but is the answer still the same ?